OpenStudy (anonymous):

Brainteasers! Medal for the first to answer correct! At a party, everyone shook hands with everybody else. There were 66 handshakes. How many people were at the party? This is just for fun, I already know the answer. First one to answer correctly gets a medal!

OpenStudy (anonymous):

let me think first

OpenStudy (anonymous):

@amistre64 @hartnn @iPwnBunnies Can you figure it out? :)

OpenStudy (cwrw238):

dont forget you cant shake your own hand

OpenStudy (anonymous):

I am working on it now

OpenStudy (ledah):

I would guess but like, I would only think of the 66 hands that were shaken, unless there is a number of people who had already shaken hands. i think i'm thinking too hard

OpenStudy (ipwnbunnies):

I'm not seeing how it can be 66, if everybody shook each other's hands...

OpenStudy (anonymous):

yea true

OpenStudy (ledah):

now that you mention it.

OpenStudy (ledah):

and then its asking how many people were at the party, and everyone has 2 hands, so if you half 66 you get 33 so i think 33 o_o

OpenStudy (anonymous):

Ok, this formula will help: \[\frac{ n(n+1) }{ 2 } = 66\] :)

OpenStudy (ledah):

LOGIC has failed me!!!!! T_T

OpenStudy (anonymous):

lol

OpenStudy (ipwnbunnies):

Are you saying people shake hands twice. ._.

OpenStudy (anonymous):

66

OpenStudy (ledah):

Hey people are weird like that.

OpenStudy (anonymous):

Think of it this way, you have 1 person who shakes everyone's hand, yet everyone does that, and in total, 66 hands were shaken, how many people were at this party :) its obviously not 66 because the total is going to be bigger than the number of people there since each person shook more than one.

OpenStudy (anonymous):

You guys want the answer?

OpenStudy (anonymous):

yes yes

OpenStudy (texaschic101):

12

OpenStudy (anonymous):

12

OpenStudy (anonymous):

Ok, the answer is 12 :)

OpenStudy (ipwnbunnies):

Ah, I see it now.

OpenStudy (anonymous):

Explanation: In general, with n+1 people, the number of handshakes is the sum of the first n consecutive numbers: 1+2+3+ ... + n. Since this sum is n(n+1)/2, we need to solve the equation n(n+1)/2 = 66. This is the quadratic equation n2+ n -132 = 0. Solving for n, we obtain 11 as the answer and deduce that there were 12 people at the party. Since 66 is a relatively small number, you can also solve this problem with a hand calculator. Add 1 + 2 = + 3 = +... etc. until the total is 66. The last number that you entered (11) is n.

OpenStudy (texaschic101):

yep :)

OpenStudy (anonymous):

Fun brainteaser site :) http://dailybrainteaser.blogspot.com/2012/12/hard-math-problems.html And don't worry, I couldn't figure it out either lol.

OpenStudy (amistre64):

assume 1 person shakes 65 hands and leaves the room then 1 person shakes 64 hands and leaves the room then 1 person shakes 63 hands and leaves the room ... when 1 person is left, the shake no hands and just leave ... how many handshakes were given?

OpenStudy (amistre64):

lol, i read the question backwards, but im sure that can be worked around to match this question

OpenStudy (anonymous):

So 66 people must have started in the room because that 1 person started shaking hands with 65 people, right?

OpenStudy (amistre64):

yes, but i read this as 66 people in a room, how many handshakes; but the sequence is arithmatic either way

OpenStudy (anonymous):

Then instead of the formula: \[\frac{ n(n+1) }{ 2 } = 66\] It will be: \[\frac{ n(n-1) }{ 2 } = 66\] right?

OpenStudy (cwrw238):

this is a good question - i'd see it before. Its pretty clever really.

OpenStudy (amistre64):

if there are n in the room, then n-1 handshakes then n-2 hand shakes then n-3 hand shakes ... then 1 handshake then 0 handshakes add up to 66 shakes altogether

OpenStudy (amistre64):

let me chk\[\frac{(n-1)(n-1+1)}{2}=66\] \[\frac{n(n-1)}{2}=66\] yes

OpenStudy (anonymous):

So then: \[n = 12\] :)

OpenStudy (amistre64):

solving the quadratic gets us n=-11 or n=12

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

And since there cannot be negative people, the answer is 12 lol